40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5879     Entry time: Sat Nov 12 02:00:36 2011
Author: Mirko 
Type: Update 
Category: Adaptive Filtering 
Subject: MC-F and other signals 

Regarding http://nodus.ligo.caltech.edu:8080/40m/5867 and http://nodus.ligo.caltech.edu:8080/40m/5869 :

MC_F signal:

The measurements on p. 5867 were done using the ADC attached to the PEM computer. There was a big difference between the MC_F signal recorded directly after the server board and the signal just before the FSS board as recorded by a PEM channel.
To understand how this happens we measured the signal at different places with a spec. analyzer:

1. WIth a locked MC measuring the signal just before the PEM ADCs (meaning after a 60ft BNC cable)
2. Same position, but unlocked and seemingly dark MC
3. Locked MC, signal just before the FSS box
4. MC_F signal that is usually going into the Pentek Generic board and is recorded in C1:IOO


=> The 60ft BNC cable adds a considerable amount of noise, but doesn't fundamentally change the signal. It is weird that the signal is white from approx. 4Hz on.
Due to Jenne's measurement ( http://nodus.ligo.caltech.edu:8080/40m/5848 ) we know the TF from MCL through PD, mixer Pentek and into C1:IOO looks like this:

This is with the double HP from 15Hz on that should be in the Pentek. So one might expect a less white signal going into the FSS board...


The dark noise in the PEM ADCs is actually a factor 10 higher than in the IOO ADCs. Still ok wrt the the seismometers.
We also tried to measure essentially the dark noise of the whole seismometer readout (seismometer box, then ADC). That seemed ok, but is of limited value since the seismometer electronics behave a bit strange when there is no seismometer connected.


Attachment 3: Compare_signals_at_all_places.fig  53 kB
Attachment 5: Channels_attached_to_the_PEM_ADC.fig  274 kB
ELOG V3.1.3-